The tip of a robot arm is moving to the right at 13m/s. Four seconds later, it is moving to the left at 2m/s. What is the magnitude of the change in velocity and what is the average acceleration? (Assume that the +x direction is the positive direction)
1
Expert's answer
2020-06-17T09:26:34-0400

We know that velocity depends on time and acceleration as follows:


"v(t)=v_0+at.\\
v(4)=-2=13+a\cdot4,\\
a=-3.75\text{ m/s}^2."

The magnitude of velocity change is


"\Delta v=|v_f-v_i|=|-2-13|=15\text{ m/s}."